• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Armstrong

    Log entry time 16:00:37 on October 23, 2009

    Entry number 296320

    keyword=Target Filled

    The target is now full, 19 K. We are on a 20 cm cell (Loop1). The target fan speed at present is 24 Hz, and we have the JT valve open at 50% for now. HP heater is at 450 W (beam off). We can open the JT valve up as required, until the usual 25 g/s or so limit. Will need to decide on fan speed, and adjust as required for the experiment.

    Input from CHL at 12 K, which is good.

    Jack and Wouter are in the Hall trying to reset Right Q2. Meanwhile, dipoles have been ramping up for a while...